OpenStudy (anonymous):

What is the arc length when Θ = 2 pi over 3 and the radius is 8 cm?

OpenStudy (anonymous):

the complete circle perimeter is \[2.8.\pi cm\] the arc lenght is \[ ((2\pi/3)/(2\pi)).16\pi = (16\pi/3) cm\]